OpenStudy (anonymous):

A curve is defined by the equation (x+y)^2-3xy=9. Find the rate of change of the slope of the curve at the point (0,3)?

OpenStudy (dumbcow):

by rate of change of slope i assume you mean the 2nd derivative differentiate and find 1st derivative \[2(x+y)(1+y') - 3y -3xy' = 0\] \[y' = \frac{y-2x}{2y-x}\] Next take 2nd derivative using quotient rule \[y'' = \frac{(y'-2)(2y-x) - (2y'-1)(y-2x)}{(2y-x)^2}\] Finally plug in given point (0,3) to obtain value for y' and then using y' , find y'' \[y' = \frac{1}{2}\] \[y'' = -\frac{1}{4}\]